OpenStudy (1018):

The perimeter of an isosceles trapezoid is 62 cm. If three sides are equal in length and the fourth side is 10 cm longer, find the area of the trapezoid.

OpenStudy (phi):

Did you draw a picture ?

OpenStudy (31356):

The perimeter of an isosceles trapezoid is 62 cm. If three sides are equal in length and the fourth side is 10 cm longer, the area of the trapezoid is given by A = (a + b)h/2 where a = (62 -- 10)/4 + 10 = 23cm, b = (62 -- 10)/4 = 13cm, h = sqrt[13^2 -- {(23 -- 13)/2}^2] = 12cm A = (23 + 13)*12/2 = 216cm^2

OpenStudy (31356):

Medal?

OpenStudy (1018):

Thsnk you! :)

OpenStudy (31356):

Your welcome